translate to a algebratic equation then solve.
the lenght of a rectangular room is six times as long as the width. the perimeter is 84 yards. find the width.

Answer by RAY100(1637)

Perimeter = 2 length +2 Width = 2(L+W)
.
But L= 6W,,,,given,,,,,substituting
.
P= 2(6W+W)=2(7W)= 14 W
.
But per = 84,,,,,,,given,,,,,substituting
.
84 = 14W
.
6=W,,,,,,answer (yards)
.
.
Check,,,, per =2L+2W= 2*6*6 +2*6=72+12=84,,,,ok
.

.